# Searching for GEMS: The Occurrence of Giant Planets orbiting M-dwarfs within 100 pc

### v1

@@ 5.4 Occurrence Rates
Given the small number of detected planets in our 100 pc sample, we instead adopt the simplified Bayesian modeling suggested in Appendix B from Hsu et al. (2018), and use a Gamma distribution prior for the occurrence rate $f_{occ,j}$, which has a shape parameter $\alpha_{0}$ and a rate parameter $\beta_{0}$. The posteriors for the occurrence rate are influenced by the choice of these parameters, especially in the regime where the number of planets is close to unity. Thus, while we use this Bayesian approach as a point of comparison for the IDEM derived occurrence for ‘Full’ M-dwarf sample listed in Table 5, and not individual spectral subtypes, we do not adopt these as the results from this work. An alternative would be to utilize a hierarchical Bayesian modeling (HBM) approach, where these parameters are dependent on the $T_{\text{eff}}$ of the sample, assuming an underlying trend. However, we leave this exercise to future work with the full 200 pc M-dwarf sample.
The mean of the gamma distribution posterior is then given by:
@@
where $n_{pl}$ is the number of planets in our full sample, and $N_{*,\rm{TESS}}$ is the effective sample size of stars. Here we note a few salient points:
  • We do not adopt a Monte Carlo approach, such as that used by Hsu et al. (2018), to estimate the expectation value of the effective stellar sample size. Instead, we use the average completeness from Figure 12, as our injection-recovery grid is sufficiently dense to provide a stable estimate without requiring resampling.
  • However, given our small planet sample concentrated below a period of 5 days, we adopt our completeness as the average of the bins in which planets or candidates exist from Figure 12. The average completeness from these bins is 7.16%, resulting in an $N_{*,\rm{TESS}}$ of $7.16\%~\times~145,612=$ 10,425 stars.
@@
  • Equation 5 does not match Equation B5 from Hsu et al. (2018), since there is a typographical error in the latter, [^25] The original equation found in Hsu et al. (2018) confines the square root to the denominator. which has been corrected here.
Following the above, we first assume $\alpha_{0}$ = $\beta_{0}$ = 1, which corresponds to an exponential distribution prior on $f_{occ}$, and get an occurrence that is comparable to that from IDEM and listed in Table 5. For illustrative purposes, we also attempt a Gaussian prior on $f_{occ}$ that matches the estimate from IDEM of $0.065^{+0.025}_{-0.027}$% by using $\alpha_{0}=6.44$, and $\beta_{0}=9763$. Doing so, we obtain a tighter posterior for $f_{occ}$. We again emphasize that given this sensitivity on prior parameters and the small sample size here, we use the IDEM derived $f_{occ}$ in subsequent discussion.
@@ 6.1 Contextualizing the occurrence of GEMS
Figure 14: Following the convention from footnote 24, we show a) occurrences of giant planets around M-dwarfs likely increases monotonically with orbital period, with close-in hot Jupiters being the rarest of all. While the total search range associated with this work is $1.0<P<10.0$ days, our occurrence rates were calculated using planets and new candidates orbiting with $1.0<P<5.0$ days. We choose to use the former range for this illustration. b) Occurrences as a function of planet radius. This figure suggests that the occurrence rate of planets orbiting M-dwarfs decreases with planet radius. No transit-based occurrence rates are available for planets with radii $4~\mathrm{R_{\oplus}}<R_{p}<8~\mathrm{R_{\oplus}}$ (Neptunes and sub-Saturns) around M-dwarfs. Limits are denoted by transparency and arrows, [^27] Ment & Charbonneau (2023) reported occurrence rates for planets with periods between 0.5 and 7.0 days, slightly different from our nominal range of 1.0–10.0 days. However, because our rates are calculated for planets with periods between 1.0 and 5.0 days, the two remain directly comparable. Similarly, Hsu et al. (2020) presented rates for periods between 0.5 and 8.0 days. and RV-based occurrence rates are denoted by triangles. In both subplots, the nominal occurrence rate found in this work is indistinguishable from the maximal value.
@@
footnote 24 shows our occurrence rates plotted against stellar mass alongside rates from previous works, spanning FGKM-dwarfs (Endl et al., 2006; Howard et al., 2012; Wright et al., 2012; Wittenmyer et al., 2020; Sabotta et al., 2021; Beleznay & Kunimoto, 2022; Mignon et al., 2025) as well as previous TESS GEMS occurrence rates from (Bryant et al., 2023; Gan et al., 2023). It is clear from this figure that the occurrence rates of giant planets across FGKM-dwarfs are consistent with a positive, monotonic relationship to stellar mass.
@@
To compare how hot-Jupiter occurrence varies with stellar mass across different surveys, we fitted the exponential relation [^28] While a fit of this form has not been explicitly produced in prior research, previous studies have employed power-law fits to model giant planet occurrence rates as a function of stellar metallicity (Johnson et al., 2010; Ghezzi et al., 2018). At the low-mass end of the stellar mass spectrum—such as that probed by our work, a power-law dependence can be reasonably approximated by an exponential model, justifying our choice of functional form. $f_{\mathrm{occ}}=a\,e^{\,b\,M_{\star}}$ to (i) the nominal occurrence rates from this work, (ii) the occurrence rates reported by Bryant et al. (2023), and (iii) those of Gan et al. (2023). All three data sets were supplemented by the occurrence rates provided by Beleznay & Kunimoto (2022) for FGK dwarfs. The best-fit coefficients, their $1\sigma$ uncertainties, and the reduced $\chi^{2}$ values are listed in Table 6. The exponential fit corresponding to this work is included, along with its $1\sigma$ region, in footnote 24.
@@
The two leading theories of giant planet formation are core accretion (Mizuno, 1980; Bodenheimer & Pollack, 1986; Pollack et al., 1996; Ikoma et al., 2000) and gravitational instability (Boss, 1997; Boss & Kanodia, 2023). In the core accretion model, material from the protoplanetary disk accumulates onto a planetesimal, eventually forming a gas giant with a solid core. In contrast, the gravitational instability model posits that portions of the disk undergo rapid collapse under their own gravity, forming gas giants without rocky cores. Both scenarios require sufficiently massive disks, making giant planet formation challenging around low-mass stars.
While the observed increase in giant planet occurrence with stellar mass alone does not decisively favor one theory, recent findings show that GEMS preferentially orbit metal-rich M-dwarfs (Gan et al., 2024; Han et al., 2024). This mirrors the planet–metallicity correlation established for FGK stars (Fischer & Valenti, 2005; Maldonado et al., 2020; Osborn & Bayliss, 2020), suggesting a shared formation pathway. Metal-rich stars host more solids necessary for core growth (Ida & Lin, 2004), strengthening the case for core accretion. Additionally, Kanodia (2024b) show a commonality in the bulk properties of transiting giant planets around FGKM dwarfs, which also hints at similar formation pathways. Together, these provide tentative support for core accretion over gravitational instability as the dominant formation pathway GEMS.
Table 6: Coefficients for the exponential fit $f_{\text{occ}}=ae^{bM_{*}}$ for the hot-Jupiter occurrence rate as a function of stellar mass.
Ref. $a$ $\sigma(a)$ $b$ $\sigma(b)$ $\chi^{2}_{\mathrm{red}}$
This work 0.0190 0.0203 3.0092 1.2486 1.69
Bryant et al. (2023) 0.1068 0.0671 1.1726 0.6627 1.29
Gan et al. (2023) 0.2574 0.2162 0.3245 0.7882 2.37
@@
14(a) shows the occurrence rates of GEMS from previous studies and from this work, plotted against orbital period (Endl et al., 2006; Sabotta et al., 2021; Mignon et al., 2025). The figure suggests a possible positive correlation between period and occurrence rate, though caution is warranted when combining results from different detection techniques. Still, such comparisons can be informative: Fernandes et al. (2019) reported a similar rise in occurrence with orbital period for FGK giant planets, with a turnover near the ice line. Identifying a comparable turnover for GEMS would be particularly interesting. While current estimates for long period GEMS remain uncertain due to smaller RV sample sizes, Gaia DR4 astrometric detections may help by probing intermediate separations (Perryman et al., 2014; Stefánsson et al., 2025b). Furthermore, observations by the Nancy Grace Roman Space Telescope (Roman; Penny et al., 2019) have the potential to detect transiting warm and cold Jupiters (Wilson et al., 2023).
@@ 6.2 Comparison with previous works
This study differs from previous work from Gan et al. (2023) and Bryant et al. (2023) in several key ways. First, the sample of M-dwarfs analyzed in this paper is the largest and includes a more comprehensive range of low-mass M-dwarfs ($M_{\star}/M_{\odot}<0.4$) compared to previous transit-based GEMS searches. We achieved this by using tglc-produced light curves, which include fainter targets than the QLP or TESS-SPOC pipelines (Table 1).
@@
Thirdly, we perform detailed spectroscopic vetting of each surviving candidate using high-resolution observations from HPF, WINERED, and NEID. This is a key improvement over previous occurrence rate studies, which typically relied on statistical false positive probabilities (FPPs) calculate exclusively with TESS photometry. Our spectroscopic follow-up was critical to our final candidate list: of the 12 new GEMS candidates identified prior, only 1 remained viable after follow-up observations. Comparing our occurrence for early M-dwarfs with those from Gan et al. (2023), and those for each of our spectral subtypes with those from Bryant et al. (2023) in footnote 24, we can see the importance of this approach to obtain an accurate estimate of giant planet occurrence, especially in the regime of faint host stars that are susceptible to background contamination and poor false positive rate estimates. We note here that close-in giant planet candidates identified in Kepler photometry also suffered from a high FP rate, although the FP rate measured in this work is considerably higher (Santerne, A. et al., 2012).
Finally, we adopt a different approach to estimate the occurrence rate and subsequent errors, which we will improve upon in our full 200 pc search. For example, Bryant et al. (2023) calculated $n_{pl}$ per bin as
@@
Bryant et al. (2023) calculated the FPP for each of their candidates using the TRICERATOPS package for a given candidate (Giacalone et al., 2021). TRICERATOPS is a statistical validation tool that uses transit photometry and stellar neighborhood information to determine the likelihood that a transit signal is caused by a false positive scenario (e.g., an eclipsing binary or background object) rather than a planet. We considered using TRICERATOPS to validate our own planet candidates; however, when testing the tool on a subset of confirmed GEMS within 200 pc using only our TGLC data, we found that it statistically validated 10 out of the 20 objects in this sample with a generous planet probability threshold of 50%. This limitation stems from the relatively large point spread function of TESS (approximately 1 arcminute) and the correspondingly large photometric uncertainties for faint M-dwarfs, which reduce the ability of TRICERATOPS to exclude false positives.
Gan et al. (2023) employed a two-step approach to estimate the FPP for each of their candidates. First, they used Forecaster (Chen & Kipping, 2017), which applies a probabilistic mass–radius relation to infer the likelihood that a transit signal arises from a low-mass star rather than a planet. Second, they estimated the probability of a candidate being a brown dwarf (BD) by dividing the number of known BDs by the number of known GEMS at the time, which does not account for detection biases, survey completeness, and the underlying occurrence rates of each population.
We follow neither approach and instead calculate the full range of possible occurrence rates given the future validation or rejection of our one new GEMS candidate.

### v3

@@ 5.4.1 Bayesian Method
We adopt the simplified Bayesian modeling method suggested in Appendix B of hsu_improving_2018 to remain statistically robust despite the small planet sample. We place a Gamma distribution prior on the occurrence rate $f_{occ}$ of each spectral subsample. The prior has a shape parameter $\alpha_{0}$ and a rate parameter $\beta_{0}$, such that the resulting posterior is the conjugate Gamma-Poisson solution for a Poisson process with mean $f_{occ}\cdot N_{*,\text{eff}}$. Because the number of planet detections is close to unity, the posterior for $f_{occ}$ is influenced by the choice of $\alpha_{0}$ and $\beta_{0}$. A more flexible alternative would be a hierarchical Bayesian modeling (HBM) approach, in which $f_{occ}$, $\alpha_{0}$, and $\beta_{0}$ vary as an assumed function of $T_{\text{eff}}$. However, we leave this exercise to future work with the full 200 pc M-dwarf sample.
The mean of the Gamma posterior for $f_{occ}$ is:
@@
where $n_{pl}$ is the number of planets in each sample, and $N_{*,\rm{TESS}}$ is the effective number of stars searched in our chosen period-radius range. We note that Equation 3 does not match Equation B5 from hsu_improving_2018, since there is a typographical error in the latter, [^22] The original equation found in hsu_improving_2018 confines the square root to the denominator. which has been corrected here.
We did not adopt a Monte Carlo approach, such as that used by hsu_improving_2018, to estimate the expectation value of the effective stellar sample size. Instead, we computed $N_{*,\text{TESS}}$ directly from our injection–recovery grid. For each spectral subsample, we calculated an effective completeness $C_{eff}$ as the arithmetic mean of the completeness values in those period–radius bins which host at least one planet (see Figure 11 and Figure 12). We then defined
$N_{*,\text{TESS}}=C_{eff}\cdot N_{*,\text{raw}}$ (4)
where $N_{*,\text{raw}}$ is the number of stars in that subsample with usable TESS data. For example, for our full sample, we obtained $C_{eff}=7.16\%$, which resulted in $N_{*,\rm{TESS}}$ of $7.16\%~\times~145,612=$ 10,425 stars.
We adopt $\alpha_{0}$ = $\beta_{0}$ = 1, which corresponds to an exponential distribution prior on $f_{occ}$. The values of $f_{occ}$ corresponding to each spectral subsample are listed in the final column of Table 5. These are the values which we quote throughout the rest of this work.
5.4.2 Inverse Detection Efficiency Method (IDEM)
For comparison, we also estimated the occurrence rate within each spectral subsample using the Inverse Detection Efficiency Method (IDEM) described by howard_planet_2012, and subsequently 2015ApJ...810...95C as well as others. Given the sparsity of planets, we did not estimate occurrences in each radius-period cell separately. Here, for the $n_{pl,j}$ planets in spectral subtype j, each planet $i$ contributes to the occurrence $f_{occ,j}$ as follows:
We accounted for errors in the period and radius estimates of all confirmed planets by bootstrapping the above equation across 5,000 iterations and accounting for the bootstrap-induced uncertainty ($\sigma_{\text{boot}}$). Our seven confirmed planets have a false positive probability (FPP) of 0. We note that bootstrap iterations in which a planet fell outside our injection bounds did not contribute to the occurrence rate.
For the uncertainty in $f_{occ,j}$, we follow recommendations from Appendix A by hsu_improving_2018, added in quadrature with errors introduced by bootstrapping:
Table 5: Occurrence rates of GEMS by spectral type
Type Median $\mathbf{M_{\star}~(M_{\odot})}$ $\mathbf{N_{\star}}$ $\mathbf{N_{\star\text{, eff}}}$ $\mathbf{N_{p}}$ $\mathbf{IDEM}$ $\mathbf{f_{\rm occ}}$ $\mathbf{\Gamma(1,1)}*$ $\mathbf{f_{\rm occ}}$    [%]
Early 0.54 27,995 3,149 2 $0.067\pm 0.047$ $0.118\pm 0.068$
Mid 0.35 37,948 2,903 4 $0.139\pm 0.069$ $0.153\pm 0.069$
Late 0.18 79,669 2,868 1 $032\pm 0.032 0.0$$036\pm 0.024$
Full 0.25 145,612 9,902 7 $065^{+0.025}_{-0.027} 0.0$$068\pm 0.024$
Note. — Each row lists the number of stars analyzed ($N_{\star}$), the effective number of stars ($N_{\star\text{, eff}}$), and the number of planets ($N_{p}$) per mass bin. Occurrence rates are computed with a simplified Bayesian approach with a $\Gamma(\alpha_{0},\beta_{0})$ prior (see Section 5.4) and, for comparison, with the Inverse Detection Efficiency Method (IDEM). Stars lacking usable TGLC data are excluded from $N_{\star}$. IDEM values are not adopted as the final results in this work, but are presented as a comparison to the hierarchical Bayesian method.
Note. — * - Adopted
Figure 13: The occurrence rates of close-in giant planets around FGKM dwarfs based on RV searches (triangles), [^24] The inclusion of RV-based occurrence rates in this figure served purely as a qualitative comparison. We leave quantitative comparisons of RV- and transit-based occurrence rates to future work. transit searches (circles), and this work (filled crosses). Upper limits are denoted by downward arrows and slight transparency, as in the case of endl_exploring_2006, howard_planet_2012, and mignon2025. The upper limit from endl_exploring_2006 is shown at a representative stellar mass of 0.5 $M_{\odot}$, whereas those for wright_frequency_2012, howard_planet_2012, and wittenmyer_cool_2020 are shown at 1 $M_{\odot}$ given that the latter consist of a sample of FGK dwarfs. Finally, the dashed black line and associated grey error envelope represent an exponential fit to the black filled crosses from this work and the purple points from beleznay_exploring_2022. Fitting to these works allows us to span the stellar mass axis using only transit photometry based studies in TESS data within a similar period and radius range to our own work.
We note here that the contribution to the occurrence rate error from bootstrapping made up $1\%$ of the total error for early- and late-type bins, due to the small number of planets occurring on the edge of a given completeness bin (see Table 5). Values of $f_{occ}$ calculated using this method are listed in Table 5.
While IDEM is computationally and intuitively simple, it has limitations, especially in the marginal detection regime where the transit S/N is comparable to the detection threshold, (i.e., for small planets, as has been highlighted by hsu_improving_2018). However, for large planets such as those under consideration here, the transit S/N is typically much larger than the detection threshold, as evidenced by our radius uncertainties in Table 3. Here, the biases induced by IDEM compared to other occurrence rate (and uncertainty) estimation techniques such as Approximate Bayesian Computing (ABC; see Figure 6 in hsu_improving_2018) are expected to be smaller.
@@ 6.1 Contextualizing the occurrence of GEMS
Figure 14: Following the convention from footnote 24, we show a) occurrences of giant planets around M-dwarfs likely increases monotonically with orbital period, with close-in hot Jupiters being the rarest of all. While the total search range associated with this work is $1.0<P<10.0$ days, our occurrence rates were calculated using planets orbiting with $1.0<P<5.0$ days. We choose to use the former range for this illustration. b) Occurrences as a function of planet radius. This figure suggests that the occurrence rate of planets orbiting M-dwarfs decreases with planet radius. No transit-based occurrence rates are available for planets with radii $4~\mathrm{R_{\oplus}}<R_{p}<8~\mathrm{R_{\oplus}}$ (Neptunes and sub-Saturns) around M-dwarfs. Limits are denoted by transparency and arrows, [^26] Ment2023 reported occurrence rates for planets with periods between 0.5 and 7.0 days, slightly different from our nominal range of 1.0–10.0 days. However, because our rates are calculated for planets with periods between 1.0 and 5.0 days, the two remain directly comparable. Similarly, hsu_occurrence_2020 presented rates for periods between 0.5 and 8.0 days. and RV-based occurrence rates are denoted by triangles. In both subplots, the nominal occurrence rate found in this work is indistinguishable from the maximal value.
@@
footnote 24 shows our occurrence rates plotted against stellar mass alongside rates from previous works, spanning FGKM-dwarfs (endl_exploring_2006; howard_planet_2012; wright_frequency_2012; wittenmyer_cool_2020; sabotta_carmenes_2021; beleznay_exploring_2022; mignon2025) as well as previous TESS GEMS occurrence rates from (bryant_occurrence_2023; gan_occurrence_2023). It is clear from this figure that the occurrence rates of giant planets across FGKM-dwarfs are consistent with a positive, monotonic relationship to stellar mass. We note that we have chosen the occurrence rate of $0.4\%\pm 0.1\%$ from howard_planet_2012, which is specific to planets between $8$ and $32\,R_{\oplus}$ with $P<10$ days, as it mimics our search bounds the closest.
@@
To compare how hot-Jupiter occurrence varies with stellar mass across different surveys, we fit the exponential relation $f_{\mathrm{occ}}=a\cdot 10^{b\cdot M_{\star}}$ to (i) the nominal occurrence rates from this work, (ii) the occurrence rates reported by bryant_occurrence_2023, and (iii) those of gan_occurrence_2023. All three data sets were supplemented by the occurrence rates provided by beleznay_exploring_2022 for transiting Jupiters around FGK dwarfs. We assumed a Gamma prior for the occurrence rates derived in this work and fitted Beta priors to those from the other three studies. We then drew $10,000$ pseudo-rates from each prior distribution and fitted the exponential parameters $a$ and $b$ for each iteration. The median values of the distributions of $a$ and $b$, their $1\sigma$ uncertainties, and the reduced $\chi^{2}$ value of the resultant fit are listed in Table 6. The exponential fit corresponding to this work is included, along with its $1\sigma$ region, in footnote 24.
@@
The two leading theories of giant planet formation are core accretion (Mizuno1980; bodenheimer_calculations_1986; Pollack96; ikoma_formation_2000) and gravitational instability (boss_giant_1997; boss_forming_2023). In the core accretion model, material from the protoplanetary disk accumulates onto a planetesimal, eventually forming a gas giant with a solid core. In contrast, the gravitational instability model posits that portions of the disk undergo rapid collapse under their own gravity, forming gas giants without rocky cores. Both scenarios require sufficiently massive disks, making giant planet formation challenging around low-mass stars.
While the observed increase in giant planet occurrence with stellar mass alone does not decisively favor one theory, recent findings show that GEMS preferentially orbit metal-rich M-dwarfs (gan_metallicity_2024; han_toi-5344_2024). This mirrors the planet–metallicity correlation established for FGK stars (fischer_planet-metallicity_2005; maldonado_hades_2020; Osborn2020), suggesting a shared formation pathway. Metal-rich stars host more solids necessary for core growth (ida_toward_2004_2), strengthening the case for core accretion. Additionally, kanodia_transiting_2024 show a commonality in the bulk properties of transiting giant planets around FGKM dwarfs, which also hints at similar formation pathways. Together, these provide tentative support for core accretion over gravitational instability as the dominant formation pathway GEMS.
Table 6: Coefficients for the exponential fit $f_{\text{occ}}=a\cdot 10^{b\cdot M_{*}}$ for the hot-Jupiter occurrence rate as a function of stellar mass.
Ref. $a$ $\sigma(a)$ $b$ $\sigma(b)$ $\chi^{2}_{\text{red}}$
This work 0.038 0.021 1.009 0.273 2.39
bryant_occurrence_2023 0.090 0.059 0.707 0.324 2.46
gan_occurrence_2023 0.265 0.166 0.250 0.288 8.46
@@
14(a) shows the occurrence rates of GEMS from previous studies and from this work, plotted against orbital period (endl_exploring_2006; sabotta_carmenes_2021; mignon2025). The figure suggests a possible positive correlation between period and occurrence rate, though caution is warranted when combining results from different detection techniques. Still, such comparisons can be informative: fernandes_hints_2019 reported a similar rise in occurrence with orbital period for FGK giant planets, with a turnover near the ice line. Identifying a comparable turnover for GEMS would be particularly interesting. While current estimates for long period GEMS remain uncertain due to smaller RV sample sizes, Gaia DR4 astrometric detections may help by probing intermediate separations (perryman_astrometric_2014; 2025AJ....169..107S). Furthermore, observations by the Nancy Grace Roman Space Telescope (Roman; Penny2019) have the potential to detect transiting warm and cold Jupiters (Wilson_2023).
@@ 6.2 Comparison with previous works
This study differs from previous work from gan_occurrence_2023 and bryant_occurrence_2023 in several key ways. First, the sample of M-dwarfs analyzed in this paper is the largest and includes a more comprehensive range of low-mass M-dwarfs ($M_{\star}/M_{\odot}<0.4$) compared to previous transit-based GEMS searches. We achieved this by using tglc-produced light curves, which include fainter targets than the QLP or TESS-SPOC pipelines (Table 1).
@@
Thirdly, we perform detailed spectroscopic vetting of each surviving candidate using high-resolution observations from HPF, WINERED, and NEID. This is a key improvement over previous occurrence rate studies, which typically relied on statistical false positive probabilities (FPPs) calculate exclusively with TESS photometry. Our spectroscopic follow-up was critical to our final candidate list: of the 13 new GEMS candidates identified prior, none remained viable after follow-up observations. Comparing our occurrence for early M-dwarfs with those from gan_occurrence_2023, and those for each of our spectral subtypes with those from bryant_occurrence_2023 in footnote 24, we can see the importance of this approach to obtain an accurate estimate of giant planet occurrence, especially in the regime of faint host stars that are susceptible to background contamination and poor false positive rate estimates. We note here that close-in giant planet candidates identified in Kepler photometry also suffered from a high FP rate, although the FP rate measured in this work is considerably higher (Santerne2012).
Finally, we adopt a different approach to estimate the occurrence rate and subsequent errors, which we will improve upon in our full 200 pc search. For example, bryant_occurrence_2023 calculated $n_{pl}$ per bin as
@@
bryant_occurrence_2023 calculated the FPP for each of their candidates using the TRICERATOPS package for a given candidate (giacalone_vetting_2021). TRICERATOPS is a statistical validation tool that uses transit photometry and stellar neighborhood information to determine the likelihood that a transit signal is caused by a false positive scenario (e.g., an eclipsing binary or background object) rather than a planet. We considered using TRICERATOPS to validate our own planet candidates; however, when testing the tool on a subset of confirmed GEMS within 200 pc using only our TGLC data, we found that it statistically validated 10 out of the 20 objects in this sample with a generous planet probability threshold of 50%. This limitation stems from the relatively large point spread function of TESS (approximately 1 arcminute) and the correspondingly large photometric uncertainties for faint M-dwarfs, which reduce the ability of TRICERATOPS to exclude false positives.
gan_occurrence_2023 employed a two-step approach to estimate the FPP for each of their candidates. First, they used Forecaster (chen_probabilistic_2017), which applies a probabilistic mass–radius relation to infer the likelihood that a transit signal arises from a low-mass star rather than a planet. Second, they estimated the probability of a candidate being a brown dwarf (BD) by dividing the number of known BDs by the number of known GEMS at the time, which does not account for detection biases, survey completeness, and the underlying occurrence rates of each population.
For this work, we relied entirely on spectroscopic followup to determine concretely whether a candidate was an FP or a real planet. We consider this the gold standard for calculating robust occurrence rates.